OpenStudy (anonymous):

Help!! Definate Medal!!! Which lines are perpendicular to 5x + 4y = 8? Choose all answers that are correct. 4y=-5x+4 4y= 5x + 12 5y = 4x + 15 5y = –4x – 15 My answer is 4y=-5x+4.. what do you think

OpenStudy (texaschic101):

5x + 4y = 8 4y = -5x + 8 y = -5/4x + 2 a perpendicular line will have a negative reciprocal slope....that means just flip the slope and change the sign. The slope in the original equation is -5/4...but a perpendicular line will have a slope of 4/5....see how I flipped the slope and changed the sign. So any equation with a slope of 4/5 will be perpendicular. 5y = 4x + 15 -- divide by 5 y = 4/5x + 3.....the slope is 4/5....that is the negative reciprocal slope of -5/4...so this is your perpendicular line.

OpenStudy (anonymous):

Thanks sooo much